# Geometric and Dynamic Scaling in Deep Transformers

### v1

@@ Abstract
Despite their empirical success, pushing Transformer architectures to extreme depth often leads to a paradoxical failure: representations become increasingly redundant, lose rank, and ultimately collapse. Existing explanations largely attribute this phenomenon to optimization instability or vanishing gradients, yet such accounts fail to explain why collapse persists even under modern normalization and initialization schemes. In this paper, we argue that the collapse of deep Transformers is fundamentally a geometric problem. Standard residual updates implicitly assume that feature accumulation is always beneficial, but offer no mechanism to constrain update directions or to erase outdated information. As depth increases, this leads to systematic drift off the semantic manifold and monotonic feature accumulation, causing representational degeneracy. We propose a unified geometric framework that addresses these failures through two orthogonal principles. First, manifold-constrained hyper-connections restrict residual updates to valid local tangent directions, preventing uncontrolled manifold drift. Second, deep delta learning introduces data-dependent, non-monotonic updates that enable reflection and erasure of redundant features rather than their unconditional accumulation. Together, these mechanisms decouple the direction and sign of feature updates, yielding a stable geometric evolution across depth. We term the resulting architecture the Manifold-Geometric Transformer (MGT). Our analysis predicts that enforcing geometric validity while allowing dynamic erasure is essential for avoiding rank collapse in ultra-deep networks. We outline an evaluation protocol for Transformers exceeding 100 layers to test the hypothesis that geometry, rather than depth itself, is the key limiting factor in deep representation learning.
@@ 1 Introduction
The residual connection is the cornerstone of modern Deep Learning, allowing gradients to flow through hundreds of layers. However, the standard additive update $\mathbf{x}_{l+1}=\mathbf{x}_{l}+\mathcal{F}(\mathbf{x}_{l})$ makes a strong assumption: that the optimal update is always an unconstrained vector addition in Euclidean space. Recent geometric deep learning perspectives suggest that data flows on low-dimensional non-linear manifolds embedded in high-dimensional space. An unconstrained update $\mathcal{F}(\mathbf{x}_{l})$ often points ”off-manifold,” leading to feature degradation and rank collapse in deep layers.
Simultaneously, the additive nature implies a ”write-only” memory mechanism. It is optimizationally difficult for a network to ”erase” information (i.e., learn $\mathcal{F}(\mathbf{x})\approx-\mathbf{x}$) when context shifts. This results in the ”residual accumulation” problem, where noise accumulates over depth.
@@
    Dynamic Traversal (via DDL): The step size along this tangent direction should be dynamic and reversible, allowing the model to perform ”Erasure” (moving backwards) or ”Identity” (staying still) operations explicitly.
@@ 2 Related Work
Manifold-Constrained Hyper-Connections (mHC). The recently proposed mHC framework (xie2025mhc) addresses the stability issues in width-scaling by constraining hyper-connections to specific geometric manifolds (e.g., the Birkhoff polytope). This ensures that the mixing of parallel streams preserves the identity mapping property intrinsic to ResNets. Our work adopts this perspective: we utilize mHC as a ”spatial regularizer” that projects the raw update vector $\mathbf{v}$ onto the valid tangent space $T_{\mathbf{x}}\mathcal{M}$ of the data manifold. While xie2025mhc focused on the geometry of the connection, we argue that the dynamics along this connection require a complementary mechanism to handle feature redundancy.
Deep Delta Learning (DDL). The theoretical foundation for our dynamic update law is grounded in the newly introduced Deep Delta Learning framework (zhang2026deep). zhang2026deep generalize the standard residual connection by modulating the identity shortcut with a learnable Delta Operator—a rank-1 geometric perturbation parameterized by a reflection direction $k(\mathbf{x})$ and a gating scalar $\beta(\mathbf{x})$. Crucially, they demonstrate that $\beta(\mathbf{x})$ enables the network to interpolate between identity mapping ($\beta\to 0$), orthogonal projection ($\beta\to 1$), and geometric reflection ($\beta\to 2$). This formulation provides the explicit mathematical mechanism for information erasure (zhang2026deep), allowing the network to actively discard historical features (”forgetting”).
The Geometric Synergy. While DDL provides the mechanism for erasure, applying it in unconstrained Euclidean space can lead to optimization difficulties. Conversely, mHC provides valid manifold constraints but lacks a dynamic mechanism to traverse backwards (erase) along the manifold. Our MGT architecture unifies these orthogonal contributions: we use mHC to define the valid direction $k(\mathbf{x})$ and DDL to control the step size and sign $\beta(\mathbf{x})$, achieving a stable yet flexible ”Erasure-and-Write” dynamic for ultra-deep scaling.
@@ 3.1 Architecture Overview
    Delta Dynamics (via DDL): Finally, the Deep Delta Learning (DDL) controller computes a dynamic gating scalar $\bm{\beta}$ based on the input context $\mathbf{X}_{l}$. This scalar modulates the rectified vector. Crucially, to strictly approximate a geometric reflection, we include a subtraction term proportional to the current state projection, allowing for true ”erasure” mechanics.
@@ 3.4 Deep Delta Learning (DDL): The Geometric Update Law
The Generalized Householder Update: To strictly implement the geometric capability of ”erasing” the current state information, the update rule must include a subtractive term proportional to the current state. We formulate the final update as:

### v2

@@ Abstract
Scaling Transformer architectures to extreme depth often leads to rank collapse: representations become redundant and degenerate despite modern normalization schemes. We argue this is fundamentally a geometric problem. Standard residual connections implicitly assume monotonic feature accumulation is beneficial, but provide no mechanism to constrain update directions or erase outdated information. As depth increases, this causes uncontrolled drift from the semantic manifold and representational collapse.
We propose the Manifold-Geometric Transformer (MGT), a unified framework addressing these failures through two orthogonal principles. First, manifold-constrained hyper-connections (mHC) restrict residual updates to valid tangent space directions, preventing manifold drift. Second, deep delta learning (DDL) enables data-dependent, non-monotonic updates that support feature erasure rather than unconditional accumulation. Together, mHC controls update direction while DDL controls magnitude and sign, yielding stable geometric evolution across depth.
Our theoretical analysis predicts that coupling geometric constraints with dynamic erasure is essential for scaling beyond current depth limits. We design a rigorous evaluation protocol for ultra-deep networks ($100+$ layers) to test whether geometry, not depth itself, is the fundamental bottleneck in Transformer scalability.
@@ 1 Introduction
The residual connection is the cornerstone of modern Deep Learning, allowing gradients to flow through hundreds of layers. However, the standard additive update $\mathbf{x}_{l+1}=\mathbf{x}_{l}+\mathcal{F}(\mathbf{x}_{l})$ makes a strong assumption: that the optimal update is always an unconstrained vector addition in Euclidean space. Recent geometric deep learning perspectives suggest that data flows on low-dimensional non-linear manifolds embedded in high-dimensional space. An unconstrained update $\mathcal{F}(\mathbf{x}_{l})$ often points “off-manifold,” leading to feature degradation and rank collapse in deep layers.
Simultaneously, the additive nature implies a “write-only” memory mechanism. It is optimizationally difficult for a network to “erase” information (i.e., learn $\mathcal{F}(\mathbf{x})\approx-\mathbf{x}$) when context shifts. This results in the “residual accumulation” problem, where noise accumulates over depth.
@@
    Dynamic Traversal (via DDL): The step size along this tangent direction should be dynamic and reversible, allowing the model to perform “Erasure” (moving backwards) or “Identity” (staying still) operations explicitly.
@@ 2 Related Work
Rank Collapse in Deep Networks. dong2021attention provided seminal theoretical analysis showing that pure self-attention layers lose rank doubly exponentially with depth, fundamentally limiting representational capacity. This collapse occurs even with modern normalization techniques (ba2016layer; xiong2020layer), suggesting that architectural modifications beyond optimization tricks are necessary. Recent depth-scaling efforts (wang2022deepnet; Touvron_2021_ICCV) have proposed specialized initialization and normalization schemes to train networks exceeding 100 layers, yet these approaches treat symptoms rather than addressing the geometric root cause of collapse.
Residual Connections and Depth. Since ResNets (he2016deep), residual connections have enabled training of deep networks by providing gradient highways. However, the standard formulation $\mathbf{x}_{l+1}=\mathbf{x}_{l}+\mathcal{F}(\mathbf{x}_{l})$ implicitly assumes additive updates in Euclidean space are universally beneficial. As Touvron_2021_ICCV observed, this monotonic accumulation can lead to feature redundancy in very deep networks. Our work addresses this through explicit erasure mechanisms.
Manifold-Constrained Hyper-Connections (mHC). The recently proposed mHC framework (xie2025mhc) addresses stability in width-scaling by constraining hyper-connections to geometric manifolds (e.g., the Birkhoff polytope), preserving the identity mapping property of ResNets. We adopt this perspective as a “spatial regularizer” that projects updates onto the tangent space $T_{\mathbf{x}}\mathcal{M}$ of the data manifold. While xie2025mhc focused on connection geometry, we argue that dynamics along these connections require complementary mechanisms for feature redundancy.
Deep Delta Learning (DDL). Our dynamic update law builds on the recent Deep Delta Learning framework (zhang2026deep), which generalizes residual connections via a learnable Delta Operator—a rank-1 perturbation parameterized by reflection direction $k(\mathbf{x})$ and gate $\beta(\mathbf{x})$. The key insight is that $\beta$ enables interpolation between identity ($\beta\to 0$), projection ($\beta\to 1$), and reflection ($\beta\to 2$), providing an explicit mechanism for information erasure that allows networks to actively forget outdated features.
Synergizing Geometry and Dynamics. While DDL provides erasure capability, applying it in unconstrained Euclidean space risks optimization difficulties. Conversely, mHC provides manifold constraints but lacks backward traversal (erasure) along the manifold. Our MGT architecture unifies these orthogonal contributions: mHC defines valid directions while DDL controls magnitude and sign, achieving stable “Erasure-and-Write” dynamics for ultra-deep scaling.
@@ 3.1 Architecture Overview
    Delta Dynamics (via DDL): Finally, the Deep Delta Learning (DDL) controller computes a dynamic gating scalar $\bm{\beta}$ based on the input context $\mathbf{X}_{l}$. This scalar modulates the rectified vector. Crucially, to strictly approximate a geometric reflection, we include a subtraction term proportional to the current state projection, allowing for true “erasure” mechanics.
@@ 3.4 Deep Delta Learning (DDL): The Geometric Update Law
The Generalized Householder Update: To strictly implement the geometric capability of “erasing” the current state information, the update rule must include a subtractive term proportional to the current state. We formulate the final update as: